OpenStudy (anonymous):

A tank can be filled by one pipe in four hours and by a second pipe in six hours; and when it is full, the tank can be drained by a third pipe in three hours. If the tank is empty and all three pipes are open, in how many hours will the tank be filled? 8 hours 12 hours 24 hours

OpenStudy (anonymous):

ok if the tank can be filled by one pipe (lets call it A) in 4 hours, how much tank can it fill in 1 hour?any idea?

OpenStudy (anonymous):

ok a Hint: if the tank can be filled by one pipe (lets call it A) in X hours, it'll fill 1/X part of the tank in 1 hour...do you follow?

hero (hero):

Pipe A = 4 hr Pipe B = 6 hr Pipe C = 3 hr \[\frac{abc}{ac + bc - ab}\] \[\frac{(4)(6)(3)}{(4)(3) + (6)(3) - (4)(6)} = \frac{72}{6} = 12\]